OpenStudy (anonymous):

Find the gradient of the secant joining the points P(a, f(a)) and Q(a+h, f(a+h)) ... Calculus!

OpenStudy (anonymous):

Here is the problem

OpenStudy (anonymous):

The gradient of the secant?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

What does that mean? I know what the secant line is, but not what the gradient of it is...

OpenStudy (anonymous):

Would it be the slope?

OpenStudy (anonymous):

Gradient is slope, yes

OpenStudy (anonymous):

Here is the problem, if its easier for you to see. I know the answer, but dont know how to get it...

OpenStudy (anonymous):

Well, basically you want to find: \[ \frac{\Delta y}{\Delta x} = \frac{Q_y-P_y}{Q_x-P_x} = \frac{f(x+h)-f(x)}{(x+h)-(x)} = \frac{f(x+h)-f(x)}{h} \]

OpenStudy (anonymous):

Yepp, I did that, but my answer was not what the textbook suggested..

OpenStudy (anonymous):

Can you show me your work a bit? Draw it or latex it.

OpenStudy (anonymous):

\[f(x+h) = 1/\left\{ (x+h)-1 \right\} \] So \[f'(x) = \left[ 1/\left\{ (x+h)-1 \right\} - 1/(x-1)\right] / h\] Sorry is that too hard to read?

OpenStudy (anonymous):

No that's fine.

OpenStudy (anonymous):

Now let's focus on this: \[ \frac{1}{x+h-1}-\frac{1}{x-1} \]How can we simplify it?

OpenStudy (anonymous):

Oh, do you cross multiply the denominator? I think I got it, thank you!